I think of a number, add 6 then multiply this answer by itself. My answer is 64. What number did I start with?

Answer by checkley79(3341)

(X+6)^2=64
X^2+12X+36=64
X^2+12X+36-64=0
X^2+12X-28=0
(X+14)(X-2)=0
X+14=0
X=-14 ANS
X-2=0
X=2 ANS
PROOF:
(-14+6)^2=64
-8^2=64
64=64
(2+6)^2=64
8^2=64
64=64